# Attention as an Adaptive Filter

### v1

@@ A.2 Stochastic Processes
  • $\mathbf{x}(t)\in\mathbb{R}^{N}$ is the latent state (or just state),
  • $\mathbf{z}(t_{k})\in\mathbb{R}^{P}$ is the observation at discrete times $t_{k}$,
  • $\mathbf{A}$ is the state matrix, and $\mathbf{B}$ is the input/noise matrix,
  • $\mathbf{C}$ is the measurement matrix,
  • $d\mathbf{w}(t)\sim\mathcal{N}(\mathbf{0},\mathbf{I}\,dt)$ is a standard Wiener process,
  • $\mathbf{v}(t_{k})\sim\mathcal{N}(\mathbf{0},\mathbf{R})$ is Gaussian measurement noise.
The continuous-time process noise covariance is defined as: $\mathbf{Q}:=\mathbf{B}\mathbf{B}^{\top}.$
Discrete-time form.
Sampling the SDE at equal intervals $\Delta t$ yields a linear Gaussian state-space model:
$\mathbf{x}_{i}=\mathbf{A}_{d}\mathbf{x}_{i-1}+\bm{\eta}_{i},\qquad\mathbf{z}_{i}=\mathbf{C}\mathbf{x}_{i}+\mathbf{v}_{i},$
with $\bm{\eta}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{Q}_{d})$ and $\mathbf{v}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{R})$. Here, $\mathbf{A}_{d}=e^{\mathbf{A}\Delta t}$ and $\mathbf{Q}_{d}=\int_{0}^{\Delta t}e^{\mathbf{A}\tau}\mathbf{Q}\,e^{\mathbf{A}^{\top}\tau}\,d\tau$.
This stochastic formulation provides the foundation for linear Gaussian estimation and filtering, including the classical Kalman Filter. Standard references include: (Antsaklis, 1997; Skogestad and Postlethwaite, 2005; Jazwinski, 1970).
A.3 The Kalman Filter
Filtering, or state estimation, is the task of estimating the states $\{\mathbf{x}_{1},\dots,\mathbf{x}_{n}\}$ from noisy measurements $\{\mathbf{z}_{1},\dots,\mathbf{z}_{n}\}$.
The Kalman Filter is the optimal recursive estimator for a linear system with Gaussian noise. It maintains a mean estimate $\mathbf{\hat{x}}_{k}$ and a covariance $\mathbf{V}_{k}$ that quantifies uncertainty, and updates them sequentially as new measurements arrive.
The Kalman Filter alternates between prediction steps (propagating the state forward using the system model) and update steps (correcting the prediction using measurements).
Prediction Step:
The model dynamics $\mathbf{A}$ propagate the previous state forward, and process noise $\mathbf{Q}$ increases uncertainty:
$\mathbf{\hat{x}}_{k|k-1}=\mathbf{A}\mathbf{\hat{x}}_{k-1},\qquad\mathbf{V}_{k|k-1}=\mathbf{A}\mathbf{V}_{k-1}\mathbf{A}^{\top}+\mathbf{Q}.$
Update Step:
When a new measurement $\mathbf{z}_{k}$ is available, the filter computes a Kalman gain $\mathbf{K}_{k}$ that weights the relative uncertainty in the prediction and the measurement:
$\mathbf{K}_{k}=\mathbf{V}_{k|k-1}\mathbf{C}^{\top}(\mathbf{C}\mathbf{V}_{k|k-1}\mathbf{C}^{\top}+\mathbf{R})^{-1}.$
The state and covariance are updated to reflect the new information:
$\mathbf{\hat{x}}_{k}=\mathbf{\hat{x}}_{k|k-1}+\mathbf{K}_{k}(\mathbf{z}_{k}-\mathbf{C}\mathbf{\hat{x}}_{k|k-1}),\quad\mathbf{V}_{k}=(\mathbf{I}-\mathbf{K}_{k}\mathbf{C})\mathbf{V}_{k|k-1}.$
When the measurement is very precise (small $\mathbf{R}$), the gain $\mathbf{K}_{k}$ is large, and the estimate relies more on the observation; when the model is more certain, the estimate relies more on the prediction.
We refer to $\mathbf{\hat{x}}_{k|k-1}$ and $\mathbf{V}_{k|k-1}$ as the a priori estimates of the state and covariance (i.e. those computed before accounting for the latest measurement), and $\mathbf{\hat{x}}_{k}$ and $\mathbf{V}_{k}$ as the a posteriori estimates (which include the latest measurement).
The difference between the actual measurement and the predicted measurement, $\mathbf{y}_{k}:=\mathbf{z}_{k}-\mathbf{C}\mathbf{\hat{x}}_{k|k-1},$ is called the innovation and represents the new information that the measurement brings relative to the model prediction. The difference between the measurement and the updated estimate, $\mathbf{r}_{k}:=\mathbf{z}_{k}-\mathbf{C}\mathbf{\hat{x}}_{k},$ is called the residual and quantifies the discrepancy after incorporating the measurement. See also: (Jazwinski, 1970; Anderson and Moore, 1979; Speyer, 2008)
Continuous-time Kalman Filtering.
For continuous-time linear stochastic systems, the Kalman Filter generalizes to the Kalman–Bucy filter. Given
$d\mathbf{x}(t)=\mathbf{A}(t)\mathbf{x}(t)\,dt+\mathbf{B}(t)d\mathbf{w}(t),\quad\mathbf{z}(t)=\mathbf{C}(t)\mathbf{x}(t)+\mathbf{v}(t),$
with process noise covariance $\mathbf{Q}(t)=\mathbf{B}(t)\mathbf{B}(t)^{\top}$ and measurement noise covariance $\mathbf{R}(t)$, the state estimate $\mathbf{\hat{x}}(t)$ evolves according to:
$\frac{d}{dt}\mathbf{\hat{x}}(t)=\mathbf{A}(t)\mathbf{\hat{x}}(t)+\mathbf{K}(t)\big{(}\mathbf{z}(t)-\mathbf{C}(t)\mathbf{\hat{x}}(t)\big{)},$
where the continuous-time Kalman gain $\mathbf{K}(t)$ is given by:
$\mathbf{K}(t)=\mathbf{V}(t)\mathbf{C}(t)^{\top}\mathbf{R}^{-1}(t),$
and the covariance $\mathbf{V}(t)$ satisfies the differential Riccati equation:
$\frac{d}{dt}\mathbf{V}(t)=\mathbf{A}(t)\mathbf{V}(t)+\mathbf{V}(t)\mathbf{A}(t)^{\top}-\mathbf{V}(t)\mathbf{C}(t)^{\top}\mathbf{R}^{-1}(t)\mathbf{C}(t)\mathbf{V}(t)+\mathbf{Q}(t).$
A.4 Adaptive Filtering
Adaptive filtering is the task of estimating or tracking a signal in real time when the system dynamics or noise statistics are unknown or time-varying. Unlike the Kalman Filter, which assumes known dynamics and noise covariances, adaptive filters adjust their parameters dynamically to minimize the mean squared error (MSE) between the filter output and a desired signal. They are called “adaptive” because the filter adapts to changes in the system or environment, usually in an online fashion.
Iteratively Reweighted Least Squares (IRLS) and Robust Kalman Filtering
Iteratively Reweighted Least Squares (IRLS) is a general approach for solving weighted least squares problems, particularly when robustness to outliers is desired. Given a set of residuals $\mathbf{r}_{i}$, IRLS solves:
$\mathbf{\hat{x}}=\arg\min_{\mathbf{x}}\sum_{i}w_{i}\,\|\mathbf{r}_{i}(\mathbf{x})\|_{2}^{2},$
where the weights $w_{i}$ are updated iteratively based on the current residuals, often using an influence function $\psi(\mathbf{r}_{i})$ such that: $w_{i}=\psi(\mathbf{r}_{i})/\mathbf{r}_{i}$.
A.5 The Attention Mechanism
Self-attention (Vaswani et al., 2023) models pairwise interactions between elements in a sequence of length $m$ by computing learned similarity scores between $d$-dimensional embeddings of each token. Each token is associated with a query, key, and value, stacked as $\mathbf{X}_{Q},\mathbf{X}_{K},\mathbf{X}_{V}\in\mathbb{R}^{m\times d}$ and linearly projected:
$\mathbf{Q}=\mathbf{X}_{Q}\mathbf{W}_{Q},\quad\mathbf{K}=\mathbf{X}_{K}\mathbf{W}_{K},\quad\mathbf{V}=\mathbf{X}_{V}\mathbf{W}_{V}.$
And the output is a weighted sum:
$\mathbf{Y}=\text{softmax}(\mathbf{Q}\mathbf{K}^{\top})\mathbf{V},\quad$
A row of $\mathbf{Y}$ can be written as:
$\mathbf{y}_{i}=\sum_{j=1}^{m}p_{ij}\mathbf{v}_{j},\quad p_{ij}=\text{softmax}(\mathbf{q}_{i}^{\top}\mathbf{k}_{j}),$
Attention can be viewed as a special-case of a precision-weighted average:
$\mathbf{y}_{i}=\Big{(}\sum_{j}\mathbf{\tilde{P}}_{ij}\Big{)}^{-1}\sum_{j}\mathbf{\tilde{P}}_{ij}\mathbf{x}_{j},$
where each $\mathbf{\tilde{P}}_{ij}$ is a positive definite matrix. This tensor formulation allows the model not only to scale each input differently, but also to rotate and stretch it in feature space.
For example, if we set:
$\mathbf{\tilde{P}}_{ij}=\frac{\mathbf{P}_{ij}}{1+\mathbf{r}_{ij}^{\top}\mathbf{P}_{ij}\mathbf{r}_{ij}},$
with $\mathbf{P}_{ij}$ diagonal and $\mathbf{r}_{ij}=\mathbf{q}_{i}-\mathbf{k}_{j}$, the normalization reduces to a simple row-wise form, leading to an element-wise version of attention.
The standard dot-product attention used in Transformers corresponds to:
$\mathbf{\tilde{P}}_{ij}=\text{softmax}(\mathbf{q}_{i}^{\top}\mathbf{k}_{j})\,\mathbf{I}.$
A.6 State Space Models
In discrete time, a linear time-invariant (LTI) system can be written as:
$\mathbf{x}_{k}=\mathbf{A}\mathbf{x}_{k-1}+\mathbf{B}\mathbf{u}_{k},\quad\mathbf{y}_{k}=\mathbf{C}\mathbf{x}_{k}+\mathbf{D}\mathbf{u}_{k},$
where $\mathbf{x}_{k}\in\mathbb{R}^{d}$ is the hidden state, $\mathbf{u}_{k}\in\mathbb{R}^{p}$ the input, and $\mathbf{y}_{k}\in\mathbb{R}^{q}$ the output.
The LTI structure allows explicit unrolling of the dynamics:
$\mathbf{y}_{k}=\mathbf{\bar{C}}\mathbf{\bar{A}}^{k}\mathbf{x}_{0}+\sum_{j=1}^{k}\mathbf{\bar{C}}\mathbf{\bar{A}}^{k-j}\mathbf{\bar{B}}\mathbf{u}_{j}=\sum_{j=0}^{k}\mathbf{\bar{h}}_{j}\mathbf{u}_{k-j},$
where $\mathbf{\bar{h}}_{j}=\mathbf{\bar{C}}\mathbf{\bar{A}}^{j}\mathbf{\bar{B}}$ is the impulse response. This shows that the system output can be viewed as a discrete convolution: $\mathbf{y}=\mathbf{\bar{h}}*\mathbf{u}$, which can be computed efficiently in the frequency domain via the Fourier or $Z$-transform.
Neural State Space Models
Recurrent models such as SSMs are prone to vanishing or exploding gradients. The HiPPO framework addresses this by constructing the state matrix $\mathbf{A}$ based on orthogonal polynomial projections, such as Legendre polynomials. The resulting HiPPO matrices optimally compress a continuous-time signal into a fixed-dimensional state while preserving information about the past. This ensures that the hidden state maintains a summary of the input sequence over long horizons, alleviating vanishing gradients (Gu et al., 2020).
Directly computing $\mathbf{y}_{k}$ involves repeated multiplications by $\mathbf{\bar{A}}$. The S4 model (Gu et al., 2022) exploits the $Z$-transform to convert the recurrence into a convolution in the frequency domain:
$\mathbf{Y}(z)=\sum_{k=0}^{\infty}\mathbf{y}_{k}z^{-k}=\sum_{i=0}^{\infty}\mathbf{h}_{i}z^{-i}\sum_{k=i}^{\infty}\mathbf{u}_{k-i}z^{-(k-i)}=\mathcal{\hat{K}}(z)\mathbf{U}(z),$
where the transfer function is:
$\mathcal{\hat{K}}(z)=\sum_{i=0}^{\infty}\mathbf{\bar{C}}\mathbf{\bar{A}}^{i}\mathbf{\bar{B}}z^{-i}.$
S4 further exploits the structure of $\mathbf{A}$ for computational efficiency. If $\mathbf{A}$ is diagonal, the transfer function simplifies:
$\mathcal{\hat{K}}_{\mathbf{\Lambda}}(z)=c(z)\sum_{i}\frac{\mathbf{C}_{i}\mathbf{B}_{i}}{g(z)-\mathbf{\Lambda}_{i}}=c(z)\cdot k_{z,\mathbf{\Lambda}}(\mathbf{C},\mathbf{B}),$
allowing parallel computation across state dimensions.
Diagonal Plus Low-Rank (DPLR) and Normal Plus Low-Rank (NPLR) Structures
To capture richer dynamics while retaining efficiency, S4 introduces a Diagonal Plus Low-Rank (DPLR) parameterization of the state matrix:
$\mathbf{A}=\bm{\Lambda}-\mathbf{P}\mathbf{Q}^{\dagger}.$
This generalizes to Normal Plus Low-Rank (NPLR) matrices:
$\mathbf{A}=\mathbf{U}\bm{\Lambda}\mathbf{U}^{\dagger}-\mathbf{P}\mathbf{Q}^{\dagger}=\mathbf{U}(\bm{\Lambda}-\mathbf{\tilde{P}}\mathbf{\tilde{Q}}^{\dagger})\mathbf{U}^{\dagger},$
showing that any NPLR matrix is unitarily similar to a DPLR matrix. This structure allows the HiPPO matrix to be embedded within S4 efficiently.
The matrix inversion lemma (Sherman–Morrison–Woodbury) enables efficient computation of $(\mathbf{\Lambda}-\mathbf{P}\mathbf{Q}^{\dagger})^{-1}$.
The Mamba models extend S4 by combining NPLR structure with careful numerical techniques to further improve stability and long-range memory (Gu and Dao, 2024). These models achieve performance competitive with Transformers on a variety of sequence modeling tasks.
@@ Appendix B Additional Material
B.1 Separable Approximation of the Full Precision Kernel
@@
$\mathbf{D}[i,j]=\mathcal{K}_{\mathbf{qk}}^{P}[k,|i-j|]\|\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]\|^{2}+\sum_{k}\|\bm{\Phi}_{\mathbf{qk}}[k,i]\|^{2}\Big{[}\mathcal{K}_{\mathbf{qk}}^{P}[k,|i-j|]\|\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\|^{2}$
$-2\mathrm{Re}\left(\mathbf{\bar{\hat{Z}}}_{\mathbf{q}}[k,i]\mathcal{K}_{\mathbf{qk}}^{P}[k,|i-j|]\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\right)\Big{]}$
If we could factor the precision tensor into separate $i$ and $j$ terms $\mathcal{P}^{C}[:,i,j]=c+\mathbf{f}[k,i]\cdot\mathbf{g}[k,j]$, then we could write:
$\mathbf{D}[i,j]=c\sum_{k}\big{\|}\bm{\Phi}_{\mathbf{qk}}[k,i]\big{\|}^{2}\Big{[}\big{\|}\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\big{\|}^{2}+\big{\|}\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]\big{\|}^{2}-2\mathrm{Re}\left(\mathbf{\bar{\hat{Z}}}_{\mathbf{q}}[k,i]\cdot\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\right)\Big{]}$
$+\sum_{k}\big{\|}\bm{\Phi}_{\mathbf{qk}}[k,i]\big{\|}^{2}\Big{[}\mathbf{f}[k,i]\cdot\mathbf{g}[k,j]\big{\|}\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\big{\|}^{2}+\mathbf{f}[k,i]\cdot\mathbf{g}[k,j]\big{\|}\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]\big{\|}^{2}$
$-2\mathrm{Re}\left(\mathbf{\bar{\hat{Z}}}_{\mathbf{q}}[k,i]\cdot\mathbf{f}[k,i]\cdot\mathbf{g}[k,j]\cdot\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\right)\Big{]}$
@@
$\mathbf{D}=c\Big{[}\big{(}\big{|}\bm{\Phi}_{\mathbf{qk}}\big{|}^{2}\big{)}^{\top}\big{|}\mathbf{\hat{Z}}_{\mathbf{k}}\big{|}^{2}+\sum_{k}\big{(}\big{|}\bm{\Phi}_{\mathbf{qk}}\big{|}^{2}\odot\big{|}\mathbf{\hat{Z}}_{\mathbf{q}}\big{|}^{2}\big{)}-2\mathrm{Re}\Big{(}\big{(}\big{|}\bm{\Phi}_{\mathbf{qk}}\big{|}^{2}\odot\mathbf{\hat{Z}}_{\mathbf{q}}\big{)}^{\dagger}\mathbf{\hat{Z}}_{\mathbf{k}}\Big{)}\Big{]}$
$+\big{(}\big{|}\bm{\Phi}_{\mathbf{qk}}\big{|}^{2}\odot\mathbf{f}\big{)}^{\top}\big{(}\big{|}\mathbf{\hat{Z}}_{\mathbf{k}}\big{|}^{2}\odot\mathbf{g}\big{)}+\big{(}\big{|}\bm{\Phi}_{\mathbf{qk}}\big{|}^{2}\odot\big{|}\mathbf{\hat{Z}}_{\mathbf{q}}\big{|}^{2}\odot\mathbf{f}\big{)}^{\top}\mathbf{g}-2\mathrm{Re}\Big{(}\big{(}\big{|}\bm{\Phi}_{\mathbf{qk}}\big{|}^{2}\odot\mathbf{\hat{Z}}_{\mathbf{q}}\odot\mathbf{f}\big{)}^{\dagger}\big{(}\mathbf{\hat{Z}}_{\mathbf{k}}\odot\mathbf{g}\big{)}\Big{)}$
Recall that the precision is:
$\mathcal{P}^{C}[:,i,j]=\left(\bm{\lambda}_{C}^{2}\odot\bm{\lambda}_{\Omega}\odot\frac{\mathbf{1}-e^{2\mathrm{Re}(\bm{\lambda})(t_{i}-t_{j})}}{-2\mathrm{Re}(\bm{\lambda})}+\bm{\lambda}_{\Gamma}\odot e^{2\mathrm{Re}(\bm{\lambda})(t_{i}-t_{j})}\right)^{-1}$
Let $\bm{\lambda}_{eff}=\frac{\bm{\lambda}_{C}^{2}\odot\bm{\lambda}_{\Omega}}{-2\mathrm{Re}(\bm{\lambda})}>0$. Then:
$\mathcal{P}^{C}[:,i,j]=\left(\bm{\lambda}_{eff}+(-\bm{\lambda}_{eff}+\bm{\lambda}_{\Gamma})e^{2\mathrm{Re}(\bm{\lambda})(t_{i}-t_{j})}\right)^{-1}$
$\frac{1}{\bm{\lambda}_{eff}}\left(1+\bm{\lambda}_{r}e^{2\mathrm{Re}(\bm{\lambda})\Delta t}\right)^{-1},\quad\text{where: }\,\bm{\lambda}_{r}=-1+\frac{\bm{\lambda}_{\Gamma}}{\bm{\lambda}_{eff}}>-1$

### v2

@@ A.3 Closed-form Precision Matrix Calculation for Diagonalizable Matrices
$\mathbf{V}_{C}(\Delta t_{ij})=\mathbf{C}\mathbf{V}(\Delta t_{ij})\mathbf{C}^{\dagger}+\bm{\Phi}_{\mathbf{z}}(\Delta t_{ij})\mathbf{R}\mathbf{M}^{\dagger}(\Delta t_{ij}),\quad\text{where}\quad\bm{\Phi}_{\mathbf{z}}(\Delta t_{ij})=\mathbf{C}e^{\mathbf{-A}\Delta t_{ij}}\mathbf{C}^{-1}.$
If $\mathbf{C}=\mathbf{S}\bm{\Lambda}_{C}\mathbf{S}^{-1}$, this becomes: $\bm{\Phi}_{\mathbf{z}}(\Delta t_{ij})=\mathbf{S}\bm{\Lambda}_{C}e^{-\bm{\Lambda}\Delta t_{ij}}\bm{\Lambda}_{C}^{-1}\mathbf{S}=\mathbf{S}e^{-\bm{\Lambda}\Delta t_{ij}}\mathbf{S}^{-1}$. If, in addition, $\mathbf{R}=\mathbf{S}\bm{\Lambda}_{R}\mathbf{S}^{\dagger}$, then:
$\mathbf{S}\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})\mathbf{S}^{\dagger}$
where:
$\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})=\big|\bm{\Lambda}_{C}\big|^{2}\bm{\Lambda}_{\mathbf{V}}(\Delta t_{ij})+e^{-2\mathrm{Re}(\bm{\Lambda})\Delta t_{ij}}\bm{\Lambda}_{R}$
$\big|\bm{\Lambda}_{C}\big|^{2}\bm{\Lambda}_{Q}\frac{\mathbf{I}-e^{2\mathrm{Re}(\bm{\Lambda})|\Delta t_{ij}|}}{-2\mathrm{Re}(\bm{\Lambda})}+e^{2\mathrm{Re}(\bm{\Lambda})|\Delta t_{ij}|}\bm{\Lambda}_{R}$
(Stable dynamics attenuate the effect of measurement noise.)
For the non-causal case,
So,
$\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})=\big|\bm{\Lambda}_{C}\big|^{2}\bm{\Lambda}_{\mathbf{V}}(\Delta t_{ij})+e^{-2\mathrm{Re}(\bm{\Lambda})\Delta t_{ij}}\bm{\Lambda}_{R}$
$\big|\bm{\Lambda}_{C}\big|^{2}\bm{\Lambda}_{Q}\frac{e^{-2\mathrm{Re}(\bm{\Lambda})\Delta t_{ij}}-\mathbf{I}}{-2\mathrm{Re}(\bm{\Lambda})}+e^{-2\mathrm{Re}(\bm{\Lambda})\Delta t_{ij}}\bm{\Lambda}_{R}$
$\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})=\text{sign}(\Delta t_{ij})\big|\bm{\Lambda}_{C}\big|^{2}\bm{\Lambda}_{Q}\mathcal{K}(\bm{\Lambda},\Delta t_{ij})+e^{-2\mathrm{Re}(\bm{\Lambda})\Delta t_{ij}}\bm{\Lambda}_{R},$
$\bm{\Lambda}_{\mathbf{V}}^{C}(\Delta t_{ij})=\sigma_{V}^{2}(\Delta t_{ij})\mathbf{I},\quad\text{where}\quad\sigma_{V}^{2}(\Delta t_{ij}):=\text{sign}(\Delta t_{ij})\big|\lambda_{C}\big|^{2}\sigma^{2}\frac{e^{-2\alpha\Delta t_{ij}}-1}{-2\alpha}+e^{-2\alpha\Delta t_{ij}}\eta^{2}$
$\dot{\lambda}_{vk}(t)=2\mathrm{Re}(\lambda_{i})\lambda_{vk}(t)+\lambda_{Qk}-\frac{\big|\lambda_{Ck}\big|^{2}}{\lambda_{Rk}}\lambda_{vk}(t)^{2}$
Letting $\alpha_{k}=\mathrm{Re}(\lambda_{k})$, $\beta_{k}=\sqrt{\alpha_{k}^{2}+\frac{\lambda_{Qk}}{\lambda_{Rk}}\big|\lambda_{Ck}\big|^{2}}$, $\gamma_{k}=\frac{1}{\beta_{k}}\log\left(\frac{\beta_{k}-\alpha_{k}}{\beta_{k}+\alpha_{k}}\right)$, the $k$th element of $\bm{\Lambda}_{\mathbf{V}}$ becomes:
$\lambda_{vk}(\Delta t_{ij})=\frac{\lambda_{Rk}}{\big|\lambda_{Ck}\big|^{2}}\left(\beta_{k}\tanh\left(\frac{\beta_{k}}{2}\left(\Delta t_{ij}-\gamma_{k}\right)\right)+\alpha_{k}\right)$
A.4 Variational Formulation of Robust Estimation
Suppose we have vector-valued measurements:
$\log p(\mathbf{\hat{z}}_{ij}\mid\bm{\mu}_{i})\geq\mathcal{L}(q,\bm{\mu}_{i}):=\mathbb{E}_{q(\mathbf{\Sigma}_{ij})}\big[\log p(\mathbf{\hat{z}}_{ij}\mid\bm{\mu}_{i},\mathbf{\Sigma}_{ij})\big]-\mathrm{KL}\big(q(\mathbf{\Sigma}_{ij})\,\big|\,p(\mathbf{\Sigma}_{ij})\big).$
$\mathbf{\Sigma}_{ij}^{-1}=\tau_{ij}\,\mathbf{P}^{C}_{ij},$
$\tau_{ij}\sim\mathrm{Gamma}(\alpha,\beta).$
The expected precision under the posterior is:
where:
The corresponding (marginal) likelihood $p(\mathbf{\hat{z}}_{ij}\mid\bm{\mu}_{i})$ is a Student’s $t$. This is equivalent to an M-estimator with loss:
$w_{ij}=\frac{1}{1+\mathbf{r}_{ij}^{\top}\mathbf{P}^{C}_{ij}\mathbf{r}_{ij}/\nu}.$
Using the fact that $\log(1+x)\approx x,\,\,\text{for }x\ll 1$, we can expand the rational form of $w_{ij}$:
@@ Appendix B Additional Material
B.1 Separable Approximation of the Precision Tensor
This section explores an alternative approach to address the memory bottleneck encountered in the diagonalizable (full tensor) case of Adaptive Filter Attention, which required $\mathcal{O}(m^{2}d)$ memory.
@@ B.1 Separable Approximation of the Precision Tensor
$\mathbf{D}[i,j]=\mathcal{P}^{C}[k,i,j]\cdot\big|\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]\big|^{2}+\sum_{k}\big|\bm{\Phi}_{\mathbf{qk}}[k,i]\big|^{2}\cdot\Big[\mathcal{P}^{C}[k,i,j]\cdot\big|\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\big|^{2}$
$-2\mathrm{Re}\left(\mathbf{\hat{Z}}^{*}_{\mathbf{q}}[k,i]\cdot\mathcal{P}^{C}[k,i,j]\cdot\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\right)\Big]$
Our previous approach — imposing a model with isotropic decay $\alpha$ and noise $\sigma,\eta$ — achieved $\mathcal{O}(m^{2})$ memory, thereby allowing the precision tensor $\mathcal{P}^{C}$ to be factored out of the summation over $k$. Here, we attempt to maintain the full $k$-dependence of the time-dependent precision tensor $\mathcal{P}^{C}[k,i,j]$ while still achieving $\mathcal{O}(m^{2}+md)$ memory. This is achieved by seeking a low-rank (separable) approximation of the precision kernel.
In particular, suppose we could approximate the precision tensor as a product of separable $i$ and $j$ terms:
$\mathcal{P}^{C}[k,i,j]\approx\mathbf{c}[k]+\mathbf{f}[k,i]\cdot\mathbf{g}[k,j].$
Then, we could write:
$\mathbf{D}[i,j]=\mathbf{c}[k]\sum_{k}\big|\bm{\Phi}_{\mathbf{qk}}[k,i]\big|^{2}\Big[\big|\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\big|^{2}+\big|\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]\big|^{2}-2\mathrm{Re}\left(\mathbf{\hat{Z}}^{*}_{\mathbf{q}}[k,i]\cdot\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\right)\Big]$
$+\sum_{k}\big|\bm{\Phi}_{\mathbf{qk}}[k,i]\big|^{2}\Big[\mathbf{f}[k,i]\cdot\mathbf{g}[k,j]\big|\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\big|^{2}+\mathbf{f}[k,i]\cdot\mathbf{g}[k,j]\big|\mathbf{\hat{Z}}_{\mathbf{q}}[k,i]\big|^{2}$
$-2\mathrm{Re}\left(\mathbf{\hat{Z}}^{*}_{\mathbf{q}}[k,i]\cdot\mathbf{f}[k,i]\cdot\mathbf{g}[k,j]\cdot\mathbf{\hat{Z}}_{\mathbf{k}}[k,j]\right)\Big]$
@@
$\mathbf{D}=\mathbf{c}[k]\Big[\big(\big|\bm{\Phi}_{\mathbf{qk}}\big|^{2}\big)^{\top}\big|\mathbf{\hat{Z}}_{\mathbf{k}}\big|^{2}+\sum_{k}\big(\big|\bm{\Phi}_{\mathbf{qk}}\big|^{2}\odot\big|\mathbf{\hat{Z}}_{\mathbf{q}}\big|^{2}\big)-2\mathrm{Re}\Big(\big(\big|\bm{\Phi}_{\mathbf{qk}}\big|^{2}\odot\mathbf{\hat{Z}}_{\mathbf{q}}\big)^{\dagger}\mathbf{\hat{Z}}_{\mathbf{k}}\Big)\Big]$
$+\big(\big|\bm{\Phi}_{\mathbf{qk}}\big|^{2}\odot\mathbf{f}\big)^{\top}\big(\big|\mathbf{\hat{Z}}_{\mathbf{k}}\big|^{2}\odot\mathbf{g}\big)+\big(\big|\bm{\Phi}_{\mathbf{qk}}\big|^{2}\odot\big|\mathbf{\hat{Z}}_{\mathbf{q}}\big|^{2}\odot\mathbf{f}\big)^{\top}\mathbf{g}$
$-2\mathrm{Re}\Big(\big(\big|\bm{\Phi}_{\mathbf{qk}}\big|^{2}\odot\mathbf{\hat{Z}}_{\mathbf{q}}\odot\mathbf{f}\big)^{\dagger}\big(\mathbf{\hat{Z}}_{\mathbf{k}}\odot\mathbf{g}\big)\Big)$
Recall that the precision tensor is defined by:
$\mathcal{P}^{C}[k,i,j]=\left(\lambda_{C,k}^{2}\odot\lambda_{Q,k}\odot\frac{1-e^{2\mathrm{Re}(\lambda_{k})(t_{i}-t_{j})}}{-2\mathrm{Re}(\lambda_{k})}+\lambda_{R,k}\odot e^{2\mathrm{Re}(\lambda_{k})(t_{i}-t_{j})}\right)^{-1}$
Let:
$\lambda_{\text{eff},k}=\frac{\lambda_{C,k}^{2}\,\lambda_{Q,k}}{-2\,\mathrm{Re}(\lambda_{k})}.$
Note that $\lambda_{\text{eff},k}>0$ for $\mathrm{Re}(\lambda_{k})<0$. The precision kernel is then,
$\mathcal{P}^{C}[k,i,j]=\left(\lambda_{\text{eff},k}+\big(-\lambda_{\text{eff},k}+\lambda_{R,k}\big)\,e^{2\,\mathrm{Re}(\lambda_{k})\,(t_{i}-t_{j})}\right)^{-1}$
$\frac{1}{\lambda_{\text{eff},k}}\left(1+\lambda_{r,k}\,e^{2\,\mathrm{Re}(\lambda_{k})\,(t_{i}-t_{j})}\right)^{-1},\quad\text{where }\lambda_{r,k}=-1+\frac{\lambda_{R,k}}{\lambda_{\text{eff},k}}>-1$